OpenStudy (anonymous):

Find the solution of square root of x+3+6=9 and determine if it is an extraneous solution. x = 0; extraneous x = 0; not extraneous x = 6; extraneous x = 6; not extraneous

OpenStudy (anonymous):

Is the radical over only x+3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You have to subtract 6 from both sides then square both sides to get rid of the radical giving you x+3=9 and then just subtract 3 and you get x=6. If you plug 6 in you get square root of 9=3+6=9=9 meaning that it is non extraneous since it works

OpenStudy (anonymous):

thanks again josh! (: